Hey!
I have been working on this problem for awhile now & any help would be greatly appreciated!
logbase 2(logbase 2(logbase 2(2^4^x))) = 3

Answer by scott8148(6628)

logbase 2(logbase 2(4^x)) = 3

logbase 2(logbase 2((2^2)^x)) = 3 ___ logbase 2(logbase 2(2^(2x))) = 3

logbase 2(2x) = 3

2x = 2^3